OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 Which of the following system of equations is not equal to the system of equations shown above?

OpenStudy (anonymous):

kk one sec

OpenStudy (anonymous):

28x - 12y = 16 and -6x + 12y = -3 14x - 6y = 4 and -14x + 28y = 1 -28x + 12y = -16 and 28x - 56y = 14

OpenStudy (anonymous):

those are 3 choices

OpenStudy (anonymous):

Any ideas?